I am not sure how to do this problem:
1/2+2/3=x/6 do i find the lcd and multiply that by the top?

Answer by rfer(12660)

1/2+2/3=x/6
3/6+4/6=x/6
7/6=x/6
6/1*7/6=x
42/6=x
7=x